# Uncertainty Quantification for Prior-Data Fitted Networks using Martingale Posteriors

### v1

@@ 4 Numerical experiments
To demonstrate the efficacy of our approach, we conduct various experiments. Here, we focus on conditional posterior and coverage properties, but we also provide an unconditional posterior estimation example and comparison with classical predictive resampling in the Appendix. We use the aforementioned combination of DPMMs and copula decomposition in all cases. To estimate $\rho$, we always randomly draw $1000$ data points simulated by the PFN.
@@ 4.1 TabPFN martingale property check
In order to investigate whether TabPFN possesses the martingale property, we fit TabPFN using $n=25$ data points of a Gamma distribution with scale and shape parameter 2. Following the martingale routine in Equation 1, we then predict $y_{n+k}$ for $k\in[1,1000]$ and always add the previous iteration’s sample to the dataset to refit TabPFN. In each step, we compute TabPFN’s PPD.
@@ 4.2 Convergence speed
We start by investigating the convergence speed analyzed in Section 3.3, in particular the results following Proposition 3.3. For this, we generate $n\in\{50,200,500,2000\}$ data points from a Gamma distribution with scale parameter 1 and shape parameter 2. After standardizing the data, we visually inspect how the posterior mean distribution of $y_{n+k}$ for $B=30$ instances evolves when running our MP approach for $k\in(0,200]$ steps.
Figure 2: Effect of number of generated samples $N$ on the convergence of $P_{N}$, depicted by $B=30$ different PPD mean estimates.
@@ Results
Figure 2 summarizes the results and confirms that we see convergence of the different trajectories within the given budget of $N=200$ sampling steps. While the convergence is not as fast as $\sqrt{n}$ for smaller sample sizes, we see a positive effect of decreased variance of CDF estimates for larger values of $n$. This suggests that for sufficiently large $n$, it is not necessary to draw more than $\sqrt{n}$ samples while $N=200$ is sufficient for smaller $n$. This is confirmed in all of our practical applications where convergence is observed no later than after $N=200$ steps.
4.3 Conditional regression posterior
As discussed in the previous section, our approach enables posterior estimation given new features $x$. We demonstrate this using two different regression data sets, a diffusion process ($n=200$) and the data from [15] ($n=400$). The diffusion process data is challenging as it evolves from an unimodal to a trimodal heteroscedastic Gaussian distribution with linear trends and sinusoidal changes depending on $x$. The second data set is used to evaluate how out-of-distribution (OOD) values for $x$ affect the estimated posterior. In both cases, we use $B=200$ replications.
@@
The resulting posteriors are depicted by their respective density values in Figure 3. Both results suggest that the processes themselves and their uncertainty are well captured. For the diffusion process, it becomes apparent from the testing data that our approach can provide very accurate values for the posterior density despite the relatively few training data points. The results for the data from [16], on the other hand, confirm that our approach is also able to detect the regions of OOD values and attribute higher probability to regions further away from the PFN prediction.
Figure 3: Visualization of PFN prediction and estimated Martingale Posterior (shaded area). Left: Diffusion process data. Right: OOD data as provided in [16].
4.4 Conditional quantile coverages
@@
4.5 Benchmark
As a last empirical evaluation, we compare our method against a bootstrap approach on real-world data from the UCI repository [4]. The bootstrap represents the natural comparison method as it can also determine uncertainty for estimated parameters from PFN without additional assumptions. In this benchmark, we run a conditional quantile regression ($\alpha\in\{0.5,0.9\}$) on the datasets airfoil ($n=1503$, $p=5$), boston ($n=252$, $p=15$), concrete ($n=1030$, $p=8$, [25]), diabetes ($n=442$, $p=10$, [5]), and fish ($n=908$, $p=6$). These datasets provide an ideal benchmark for PFNs, which are still limited in their scalability to larger datasets. We use $1/5$th of the dataset for training and aim to measure the uncertainty of PFN for the predicted quantile on the other $4/5$th of the dataset. In addition to our approach using the martingale posterior with $B=100$, we also use a non-parametric bootstrap by rerunning the PFN method for $B=20$ times by drawing samples from the training dataset with replacement, making quantile predictions, and compute the empirical 90% confidence interval of these predictions. We use both methods to construct a 90% credible/confidence interval (CI) and check the coverage of the predicted quantile given test dataset features. We also compute the length of these intervals to see how conservative our method is in comparison to a bootstrap approach, as well as the time for computation. We repeat the process five times with different splits.
Results
Table 2 summarizes the results of our benchmark for $\alpha=0.5$, i.e., conditional median regression. While both methods often have very close to nominal coverage or provide larger than nominal coverage, the bootstrap method fails in one case to provide a CI with close to nominal coverage. Despite having larger than nominal coverage, our MP approach consistently yields shorter or equally large CIs while often being twice as fast as the bootstrap.
@@
Martingale Posteriors ($B=100)$ Bootstrap ($B=20)$
@@
In Table 3, we further show results for the quantile regression with $\alpha=0.9$. The results are similar to those of the median regression with our approach taking a fraction of the bootstrap approach while yielding notably smaller CIs. The coverage of both methods is often close to the nominal level, while the bootstrap again struggles to estimate intervals close to nominal level for the concrete dataset.
@@
Martingale Posteriors ($B=100)$ Bootstrap ($B=20)$
@@
5 Discussion
This work proposes an efficient and principled Bayesian uncertainty quantification method for estimates derived from prior fitted networks.
Limitations
Posteriors for other in-context learners
Our theory and algorithm can be transferred directly to other in-context predictors that provide a PPD; for example, when using large language models (LLMs) as a prediction tool for numeric value prediction, such as in [6], or by generating so-called LLM processes [21]. Both approaches use a text-based prompt to predict future numeric values, thereby reducing the large token space to, e.g., $[0,1]$ [6] or a 10-dimensional digit space [21].
We provide a proof-of-concept by prompting different LLMs to construct a CDF based on $n=25$ using the setup of [6]. The prompt is given in Section B.4. Using the LLM output, we then run our MP approach for $N=1000$ sampling iterations. The results are depicted in Figure 4, showing that the martingale property holds for our approach (the expected value of our generated blue CDFs coincides with the initial CDF of the LLM). While the blue curves also nicely depict the epistemic uncertainty of the LLM for the predicted CDFs, it becomes clear that the provided method—as expected—cannot counteract biases in the LLM (as is the case for GPT 4o). For more advanced LLMs, however, our approach gives uncertainty quantification akin to what is investigated [21], but potentially much cheaper.
Posteriors for LLM-generated text
PFNs are in-context learners based on the transformer architecture, just as modern LLMs. LLMs are trained to approximate the PPD $p(y|x,\mathcal{D})$ of the next token $y$ given a context $x$ of previous tokens and the corpus $\mathcal{D}$ they were trained on. Given this similarity, the martingale posterior approach could potentially also be adapted to quantify uncertainty in LLM-generated text. There are, however, two major obstacles. First, simple PPD updates as in Equation 2 are impossible due to the complex relation between the next and the previous tokens. The complexity of such updates can likely only be matched by an equally strong LLM. Second, Proposition 3.1 and Proposition 3.2 fail due to the lack of exchangeability in the generated output tokens. Contraction results like Proposition 3.3, however, rely on only the martingale property and would remain valid, although the interpretation of the distributions $P_{M}$ for $M\to\infty$ is less clear. Investigating how and to what extent martingale posterior ideas transfer to language models is an exciting path for future research.
Appendix A Proof of Proposition 3.3
We use arguments similar to [8]. It holds
$\mathbb{E}[P_{M}(y)|y_{(n+1):(n+M-1)}]=(1-\alpha_{n+M})P_{M}(y)+\alpha_{n+M}\mathbb{E}_{y_{n+M}\sim P_{M-1}}[H_{\rho}(P_{M-1}(y),P_{M-1}(y_{n+M})].$
by the properties of the Gaussian copula. Hence,
$\mathbb{E}[P_{M}(y)|y_{(n+1):(n+M-1)}]=P_{M-1}(y),$
which implies that $P_{M}(y)$ is a martingale. Furthermore,
$|P_{M}(y)-P_{M-1}(y)|\leq\alpha_{n+M}\leq\frac{2}{n+M+1},\quad\text{for all }y\in\mathbb{R},$
and
[…]
Now the Azuma-Hoeffding inequality (e.g. [1]) yields
$\Pr(|P_{M}(y)-P_{N}(y)|\geq\epsilon)\leq 2\exp\left(-\frac{\epsilon^{2}}{2\sum_{i=N}^{\infty}\alpha_{n+i}^{2}}\right)\leq 2\exp\left(-\frac{\epsilon^{2}(n+N)}{8}\right).$
Appendix B Further numerical experiments and details
B.1 Unconditional quantile posterior
We here provide another experiment where we analyze the ability of our approach to estimate the posterior of an extreme quantile of a skewed distribution. For this, we simulate a Gamma distribution with different shape parameters $\gamma\in\{1,2,4\}$, inducing varying left-skewness. We then task PFN to estimate the 99%-quantile (the function $T$) and use our approach to compute the posterior uncertainty for PFN’s estimate. In this experiment, we use $B=1,000$ replications and $N=10,000$. To make the task of quantile estimation more challenging, we use a relatively small data set size of $n=25$. To evaluate the performance, we compare the distribution against the true value and a posterior estimate by the predictive resampling approach from [8] that does not have access to the PFN.
Results
B.2 Data generation for diffusion process

### v2

@@ 4 NUMERICAL EXPERIMENTS
To demonstrate the efficacy of our approach, we conduct various experiments. Here, we focus on conditional posterior and coverage properties, but we also provide further ablation studies to better understand our method’s behaviour. We use the aforementioned combination of DPMMs and copula decomposition in all cases. To estimate $\rho$, we always randomly draw $1000$ data points simulated by the PFN. We use the original learning rate approach in the first three sections, but perform dedicated ablations for benchmarks. To this end, we will refer to our method with the learning rate proposed in Fong et al., (2023) as Ours (default), while the adaptations proposed in Section 3.3 based on theoretical considerations and following the discussion about the biasedness of TabPFN are called Ours (type 1) and Ours (type 2).
@@ 4.1 TabPFN Martingale Property Check
In order to investigate whether TabPFN possesses the martingale property, we fit TabPFN using $n=25$ data points of a Gamma distribution with scale and shape parameter 2. Following the martingale routine in (1), we then predict $y_{n+k}$ for $k\in[1,1000]$ and always add the previous iteration’s sample to the dataset to refit TabPFN. In each step, we compute TabPFN’s PPD.
@@ 4.2 Convergence Speed
Figure 3: Effect of NN on the convergence of $P_{N}$, depicted by $B=30$ different PPD mean estimates.
We start by investigating the convergence speed of the MP sampling algorithm. For this, we subsample the concrete dataset (Yeh,, 1998) to generate datasets of different sizes. After standardizing the $n\in\{25,100,250,1000\}$ data points, we visually inspect how the conditional posterior mean distribution of $y_{n+k}$ conditional on fixed features $x$ for $B=30$ instances evolves when running our MP approach for $k\in(0,1000]$ steps. Figure 3 summarizes the results for the default type learning rate and confirms that the different trajectories stabilize within the given budget of $N=1000$ sampling steps independent of $n$. This is confirmed in all of our practical applications, where convergence is observed no later than after $N=500$ steps. Results for different learning rate schedules are given in Section B.2.3.
4.3 Conditional Regression Posterior
Our approach enables posterior estimation given new features $x$, complementing PFN’s aleatoric uncertainty with epistemic uncertainty. We demonstrate this using a diffusion process ($n=200$). This data is challenging as it evolves from an unimodal to a trimodal heteroscedastic Gaussian distribution with linear trends and sinusoidal changes depending on $x$. We run our routine and compute both the estimated density by TabPFN and the variance of this density as a measure of epistemic uncertainty provided by our method.
@@ Results
The results are depicted in Figure 1. While TabPFN is able to fit the data-generating process well and its uncertainty (normalized density values) aligns with the apparent data for larger $x$ values, higher density values also coincide with higher uncertainty (as measured by the normalized density variance). This becomes particularly visible in the lower branch of the diffusion process, where less training data is available. The provided density by TabPFN is comparatively small in this area, while the corresponding epistemic uncertainty is relatively large.
4.4 Benchmark
Figure 4: Mean values and standard deviations (error bars) of miscoverage (observed minus nominal value; x-axis) and interval length (y-axis) of the 90% credible intervals produced by the original method (blue), our proposed adaptions (orange, purple), and the bootstrap, across the different benchmark datasets. The red vertical line indicates perfect coverage.
As an empirical evaluation, we compare our method using the previously presented learning rate approaches on real-world data from the UCI repository (Dua and Graff,, 2017). We further compare against a bootstrap approach, which represents the natural comparison method, as it can also determine uncertainty for estimated parameters from PFN without additional parametric assumptions. In this benchmark, we run a conditional quantile regression ($\alpha\in\{0.5,0.9\}$) on the datasets airfoil ($n=1503$, $p=5$), boston ($n=252$, $p=15$), concrete ($n=1030$, $p=8$, Yeh,, 1998), diabetes ($n=442$, $p=10$, Efron et al.,, 2004), energy ($n=768$, $p=8$, Tsanas and Xifara,, 2012), and fish ($n=908$, $p=6$). These datasets provide an ideal benchmark for PFNs, which are still limited in their scalability to larger datasets. We use $n=100$ data points for training and aim to measure the uncertainty of PFN for the predicted quantile on the remaining data points. In addition to our approaches using the martingale posterior with $B=100$, the non-parametric bootstrap is computed. The bootstrap reruns the PFN method for $R=20$ times by drawing samples from the training dataset with replacement, makes quantile predictions, and computes the empirical 90% confidence interval of these predictions. We use both methods to construct a 90% credible/confidence interval (CI) and check the coverage of the predicted quantile given test dataset features. We also compute the length of these intervals to see how conservative our method is in comparison to a bootstrap approach, as well as the time for computation. We repeat the process five times with different splits.
@@
Figure 4 summarizes the results of our benchmark for $\alpha=0.5$, i.e., conditional median regression. The results are similar to those with smaller dataset size and for the quantile regression case (provided in Tables 3 and 4 in Appendix B), with our approach taking a fraction of the bootstrap approach’s time while yielding notably smaller CIs with better coverage. When comparing the different learning rate schemes, we see that in most cases, it is possible to either decrease the interval width (airfoil, boston, concrete, energy, fish) or improve the coverage (boston, concrete, energy, fish), or both, over the approach proposed in Fong et al., (2023).
Runtime
In addition, Section B.2.5 provides a comparison that fixes the time budget of both the MP and the bootstrap approach. As the bootstrap takes notably longer than our method (requiring $R$ forward passes of the PFN while our routine only uses PFNs once for $P_{0}$ and the estimation of $\rho$), the undercoverage of the bootstrap becomes even more pronounced.
4.5 Ablation: Dataset Size
In the previous sections, the sizes of the datasets are relatively small. Using small datasets is done intentionally, as PFNs particularly shine in small-scale applications by leveraging their pretraining. Here, we further investigate how coverages change when increasing the dataset size using the two largest datasets (airfoil, concrete) from Section 4.4 and subsample the datasets to training data sizes of $n\in\{50,100,200,500,700\}$.
Failure due to Bias
This analysis is particularly interesting in light of our discussion in Section 3.5, suggesting that intervals will not yield nominal coverage, even for $n\to\infty$ due to the non-vanishing bias of TabPFN. Empirically, this is validated in Figure 5, showing that neither the learning rate from Fong et al., (2023) nor our suggested alternative (type 1) obtains nominal coverage and even drifts away from the nominal level for increasing $n$.
Figure 5: Coverage and interval length behavior for the larger UCI datasets (rows) for differently sized training datasets.
In contrast, when using the type 2 approach suggested in Section 3.3, we find that our method is working as expected with decreasing interval lengths and decreasing deviations from the nominal coverage (0.9) for an increase in $n$.
5 Discussion
This work proposes an efficient and principled Bayesian uncertainty quantification method for estimates derived from prior fitted networks. Our method thereby provides epistemic uncertainty complementing the aleatoric uncertainty of PFNs. Epistemic uncertainty methods such as deep ensembles (Lakshminarayanan et al.,, 2017) are a potential alternative, but computationally expensive, as they require training multiple models. In the context of pretrained models, this is usually not a feasible option. In contrast, our method only retrains the martingale posterior routine, not the model itself, which is generally much more efficient.
Limitations
Posteriors for Other In-Context Learners
Our theory and algorithm can be transferred directly to other in-context predictors that provide a PPD; for example, when using large language models (LLMs) as a prediction tool for numeric value prediction, such as in Falck et al., (2024), or by generating so-called LLM processes (Requeima et al.,, 2024). Both approaches use a text-based prompt to predict future numeric values, thereby reducing the large token space to, e.g., $[0,1]$ (Falck et al.,, 2024) or a 10-dimensional digit space (Requeima et al.,, 2024).
LLM Proof-of-Concept
To substantiate our previous hypothesis, we provide a proof-of-concept by prompting different LLMs to construct a CDF based on $n=25$ using the setup of Falck et al., (2024). The prompt is given in Section B.2.8. While Falck et al., (2024) investigated whether different LLMs possess the martingale property, we are interested in whether our approach can provide uncertainty for these models. To do so, we simulate i.i.d. data from a Gaussian distribution and provide it to the LLM (GPT-4o, GPT-o3-mini-high, DeepSeek R1) including a context that specifies how the data was generated. We then ask the LLM to generate CDF values $F(y_{i})$ for this distribution for a prespecified grid of $y_{i}$ values. Figure 6 shows the initial output CDF of the LLM as a red dashed line. The underlying true CDF is marked with a dashed-dotted green line, indicating that GPT-4o is unable to predict the CDF, while the others do so correctly. We then run $N=1000$ replications of forward loops of our Algorithm 1 to sample values $y_{(n+1):(n+N)}^{(b)}$ and plot the resulting CDF lines for $y_{1:(n+N)}^{(b)},b=1,\ldots,B$ in blue. These lines indicate the uncertainty of the CDF generated by the LLM. The average CDF resulting from these blue curves is plotted in yellow, indicating that all models have the martingale property (although GPT-4o is not correct in its prediction).
While the blue curves also nicely depict the epistemic uncertainty of the LLM for the predicted CDFs, it becomes clear that the provided method—as expected—cannot counteract biases in the LLM (as is the case for GPT 4o). For more advanced LLMs, however, our approach gives uncertainty quantification akin to what is investigated Requeima et al., (2024), but potentially much cheaper.
Posteriors for LLM-Generated Text
PFNs are in-context learners based on the transformer architecture, just as modern LLMs. LLMs are trained to approximate the PPD $p(y|x,\mathcal{D})$ of the next token $y$ given a context $x$ of previous tokens and the corpus $\mathcal{D}$ they were trained on. Given this similarity, the martingale posterior approach could potentially also be adapted to quantify uncertainty in LLM-generated text. There are, however, two major obstacles. First, simple PPD updates, as in (3), are impossible due to the complex relation between the next and the previous tokens. The complexity of such updates can likely only be matched by an equally strong LLM. Second, Proposition 1 and Proposition 2 fail due to the lack of exchangeability in the generated output tokens. Contraction results like Proposition 3, however, rely on only the martingale property and would remain valid, although the interpretation of the distributions $P_{M}$ for $M\to\infty$ is less clear. Investigating how and to what extent martingale posterior ideas transfer to language models is an exciting path for future research.
Supplementary Materials
Appendix A PROOF OF PROPOSITION 3
We first show that $P_{M}(y)$ is a martingale. It holds
$\mathbb{E}[P_{M}(y)|y_{(n+1):(n+M-1)}]=(1-\alpha_{n+M})P_{M-1}(y)+\alpha_{n+M}\mathbb{E}_{y_{n+M}\sim P_{M-1}}[H_{\rho}(P_{M-1}(y),P_{M-1}(y_{n+M})].$
by the properties of the Gaussian copula. Hence,
$\mathbb{E}[P_{M}(y)|y_{(n+1):(n+M-1)}]=P_{M-1}(y),$
which implies that $P_{M}(y)$ is a martingale adapted to the filtration $\sigma(y_{1:n})_{n\in\mathbb{N}}$.
Next, we apply the uniform-in-time version of the Hoeffding-Azuma inequality given in Corollary 1 (a) of Howard et al., (2020). It holds
$|P_{M}(y)-P_{M-1}(y)|\leq\alpha_{n+M}\leq C(n+M+1)^{-\beta},\quad\text{for all }y\in\mathbb{R},$
and, for all $M\geq N$,
$\sum_{i=N}^{M}\alpha_{n+i}^{2}\leq\sum_{i=N}^{\infty}\alpha_{n+i}^{2}\leq C^{2}\int_{N}^{\infty}(n+t)^{-2\beta}dt\leq C_{\beta}(n+N)^{1-2\beta},$
where $C_{\beta}=C^{2}/(2\beta-1)$.
Now Corollary 1 (a) of Howard et al., (2020) with $m=C_{\beta}(n+N)^{1-2\beta})$ yields
$\Pr(\exists M\geq N\colon|P_{M}(y)-P_{N}(y)|\geq\epsilon)\leq 2\exp\left(-\frac{\epsilon^{2}}{2m}\right)=2\exp\left(-\frac{\epsilon^{2}(n+N)^{2\beta-1}}{2C_{\beta}}\right).$
Setting ${\epsilon}=\sqrt{C_{\beta}\log(2/\delta)}(n+N)^{1/2-\beta}$ gives,
$\Pr(\exists M\geq N\colon|P_{M}(y)-P_{N}(y)|\geq\epsilon)\leq\delta.$
Now the claim follows since $P_{M}(y)\to P_{\infty,x}(y)$ almost surely (Proposition 2). ∎
Appendix B FURTHER NUMERICAL EXPERIMENTS AND DETAILS
B.1 Estimation of $\rho$
We implemented the estimation of $\rho$ for the Gaussian copula update as follows: If the computational budget is limited, we follow the routine suggested by Fong et al., (2023) and maximize the prequential log score $\sum_{i=1}^{n}\log p_{i-1}(y_{i}|\bm{x}_{i})$ for $\rho$ (an efficient method is implemented in the MP module by the same authors). If the dataset size is limited or if one likes to separate tuning from training, one could potentially draw a new and larger set of density values from TabPFN and independently estimate using this “tuning set”. In practice, we found the training data to be sufficient for good estimation of $\rho$. We also reuse this estimator for all predictive steps in our Martingale Posterior computation and have not found any significant improvement by re-estimating the correlation or manually tuning via cross-validation.
B.2 Additional Experiments
B.2.1 Unconditional Quantile Posterior
We here provide another experiment where we analyze the ability of our approach to estimate the posterior of an extreme quantile of a skewed distribution. For this, we simulate a Gamma distribution with different shape parameters $\gamma\in\{1,2,4\}$, inducing varying left-skewness. We then task PFN to estimate the 99%-quantile (the function $T$) and use our approach to compute the posterior uncertainty for PFN’s estimate. In this experiment, we use $B=1,000$ replications and $N=10,000$. To make the task of quantile estimation more challenging, we use a relatively small dataset size of $n=25$. To evaluate the performance, we compare the distribution against the true value and a posterior estimate by the predictive resampling approach from Fong et al., (2023) that does not have access to the PFN.
Results
B.2.2 Conditional Quantile Coverages
@@
Boostrap Comparison
To show that this is a difficult task, we further provide results for the bootstrap in Table 2. The results show that the bootstrap (with $B=400$) also struggles to obtain nominal coverages, and for all larger x- and more extreme quantile values, with significantly lower coverages than 90%.
$x\backslash\alpha$ 0.05 0.10 0.25 0.50 0.75 0.90 0.95
0.2 0.65 $\pm$ 0.30 0.70 $\pm$ 0.29 0.70 $\pm$ 0.29 0.80 $\pm$ 0.25 0.80 $\pm$ 0.25 0.70 $\pm$ 0.29 0.65 $\pm$ 0.30
0.5 0.70 $\pm$ 0.29 0.85 $\pm$ 0.26 0.65 $\pm$ 0.30 0.75 $\pm$ 0.28 0.80 $\pm$ 0.25 0.70 $\pm$ 0.29 0.65 $\pm$ 0.30
0.8 0.40 $\pm$ 0.31 0.55 $\pm$ 0.32 0.75 $\pm$ 0.28 0.80 $\pm$ 0.25 0.70 $\pm$ 0.29 0.50 $\pm$ 0.32 0.60 $\pm$ 0.31
Table 2: Results for different values of $x$ and $\alpha$.
B.2.3 Convergence Speed: Additional Results
Figures 8 and 9 depict the convergence speeds for the analysis in Section 4.2 for the two other learning rate schedules, type 1 and type 2, respectively. In this example, all results look qualitatively very similar.
Figure 8: Effect of number of generated samples NN on the convergence of $P_{N}$, depicted by $B=30$ different PPD mean estimates.
Figure 9: Effect of number of generated samples NN on the convergence of $P_{N}$, depicted by $B=30$ different PPD mean estimates.
B.2.4 Benchmarks: Additional Results
Tables 3 and 4 summarize the results of our benchmark for $\alpha=0.5$ and $\alpha=0.9$, respectively for $n=25$. While both methods often have very close to nominal coverage or provide larger than nominal coverage, the bootstrap method fails in one case to provide a CI with close to nominal coverage. Despite having larger than nominal coverage, our MP approach consistently yields shorter or equally large CIs while often being twice as fast as the bootstrap.
@@ B.2.4 Benchmarks: Additional Results
Martingale Posteriors ($B=100)$ Bootstrap ($R=20)$
@@
Martingale Posteriors ($B=100)$ Bootstrap ($R=20)$
@@
B.2.5 Benchmarks: Comparison Using Fixed Time Budget
To make a fair comparison between our method and the bootstrap, we run additional benchmarks with an equal time budget for both methods. The results are given in Table 5. These results further strengthen the case for our method by showing that for equal time budget, the Bootstrap produces much wider intervals that nevertheless fail to reach the target coverage in several cases.
Dataset Coverage (MP) Interval Length (MP) Coverage (Bootstrap) Interval Length (Bootstrap)
airfoil 0.96 $\pm$ 0.01 0.23 $\pm$ 0.05 0.87 $\pm$ 0.03 0.40 $\pm$ 0.04
boston 0.90 $\pm$ 0.01 0.25 $\pm$ 0.02 0.92 $\pm$ 0.06 0.42 $\pm$ 0.03
concrete 0.91 $\pm$ 0.01 0.33 $\pm$ 0.04 0.64 $\pm$ 0.03 0.38 $\pm$ 0.03
diabetes 0.90 $\pm$ 0.03 0.65 $\pm$ 0.16 0.82 $\pm$ 0.07 0.82 $\pm$ 0.13
energy 0.97 $\pm$ 0.01 0.09 $\pm$ 0.01 0.73 $\pm$ 0.10 0.08 $\pm$ 0.01
fish 0.92 $\pm$ 0.02 0.38 $\pm$ 0.05 0.91 $\pm$ 0.05 0.64 $\pm$ 0.06
Table 5: Coverage and interval lengths for MP and Bootstrap methods.
B.2.6 Data Generation for Diffusion Process